Find an equation where the y – intercept = 4 and x intercept = -4.
Answer by mananth(16946)

the points of intercept are (0,4) (-4,0)
x1 y1 x2 y2
0 4 -4 0

slope m = (y2-y1)/(x2-x1)
( 0 - 4 )/( -4 - 0 )
( -4 / -4 )
m= 1

Plug value of the slope and point (0,4) in
Y = m x + b
4.00 = 0 + b
b= 4

So the equation will be
Y =x +4
m.ananth@hotmail.ca